bond distance C-F in (CF4) & Si-F in (SiF4) are respective 1.22 A* & 1.54 A*. C-Si bond is 1.87 A*. Calculate the covalent radius of F atom ignoring the electronegativity differences. 

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
bond distance C-F in (CF4) & Si-F in (SiF4) are respective 1.22 A* & 1.54 A*. C-Si bond is 1.87 A*
let the radius of carbon,silicon and fluorine atom be x,y and z respectively
now the sum of the radius of each atom will be equal to the bond distance
therefore
x +y =1.87
y+z=1.54
z+a =1.22
now solving these three equation we get the radius of fluorine
as .125A